I have two problems I need help with. The first one is {8x+1} = x+2 and the second one is 4{2x} + 4{x+3} = 0 We tried several times not being able to solve. Thank you in advance.
Answer by queenofit(120)

If you have your ( ) in the correct places this is what I got for your first equation.
8x+1=x+2 you must move the variables to one side and the numbers to the other
8x-x=2-1 remember to change your signs when you cross the = sign
7x=1 then divide each side by 7 because you want your answer to be x=
x= 1/7

equation 2
4(2x)+ 4(x+3)=0 first you must distribute the number oustide the ( ) to the inside of the ( ) by multiplication as followes.
(4*2x)+(4*x)+(4*3)= 0 this gives you the following
8x+4x+12=0 combine the like terms
12x+12=0 move the 12 to the other side of the = sign be subtracting 12 from each side as you did in the first equation you asked about.
12x+12-12=0-12
12x=-12 divide each side by 12 to get your x alone
12x/12= -12/12
x= -1
I hope this helps, if these are not the correct answers then there is a problem with where your ( ) are. Remember, the answers in the back of your textbook are not always correct. grad. students get paid to do those and sometimes they are in a hurry and they get them wrong.